Fermi National Laboratory
 

Operation Summary for January 26, 2011 - January 28, 2011

 

Wednesday Jan. 26

The day shift began with Operations monitoring the Tevatron (TEV) store 8446 (97.6E30) and the Recycler stash (309.2E10), with the Antiproton Source (Pbar) stacking (22.9E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, with the Meson FTBF T-978 (CALICE) experiment off due to beamline vacuum problems, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8446 at 10:47 AM, put the TEV into HEP shot setup, and established store 8447 at 12:17 PM with an average initial luminosity (AIL) of 313.81E30. Operations reported that the Tevatron quenched at TEV sector E1 due to a C49 vertical Separator spark.

During the evening shift, Operations recovered the TeV, turned on, and put the TEV through a dry squeeze at 4:13 PM. TEV personnel conducted studies through to the end of shift.

Thursday Jan. 27

The midnight shift began with Operations monitoring the stash (321.4E10), with Pbar stacking (11.0E10), with T-978 still off due to vacuum problems, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TEV into shot setup and established store 8454 at 1:23 AM with an AIL of 277.78E30.

During the day shift, at 8:23 AM, Operations reported that store 8454 quenched at 8:23 AM due to a Proton Abort Kicker (PAK3) pre-fire. Following the completion of the M-Test beamline vacuum repairs, Operations established beam to T-978 at 11:09 AM. Operators completed the TEV recovery, turn-on, and checkout at 3:19 PM. Operations put the TEV into shot setup.

During the evening shift, Operations established store 8458 at 4:17 PM with an AIL of 333.34E30.

Friday Jan. 28

The midnight shift began with Operations monitoring store 8458 (118.64E30) and the stash (233.0E10), with Pbar stacking (10.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a network storm, which started from a NuMI Front End (FE), took down CAMAC networks at 4:25 AM. Operators rebooted the FE and had most systems recovered by 6:15 AM. Booster personnel found a bad controller in their Gradient Magnet Power Supply (GMPS) while trying to recover its FE; they replaced the controller and tuned through to the end of shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/28/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory